OpenStudy (anonymous):

help please?? :c Find the range of f(x) = -x + 4 for the domain (-3, -2, -1

OpenStudy (anonymous):

Domain is all the x values, and Range is all the why values. Substitute x with the set you have. Example: f(-3) = -(-3) + 4

OpenStudy (anonymous):

*y values.

OpenStudy (anonymous):

I don't understand this at all

OpenStudy (anonymous):

f(x) = -x + 4 is the same as y = -x + 4

OpenStudy (anonymous):

the choices are 7, 6 5, 3 -7, -6, -5, 3 7, 6, 5, 4 -7,-6,-5,-4

OpenStudy (anonymous):

Example: Domain: {-3, -2, -1, 0} --- Range: {?} f(-3) = -(-3) + 4 = (3) + 4 = 7 Range for -3 is 7.

OpenStudy (anonymous):

It's best if you try it out. It'll allow you to get the grip on things :)

OpenStudy (anonymous):

f(x) = -x + 4 Keep replacing x with the value provided in the domain until you have all range.

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

-(x) = -1 (x)

OpenStudy (anonymous):

okay so I just tried.. would it be 7, 6 5, 3

OpenStudy (anonymous):

so 7, 6, 5, 4 ?

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

incorrect >.<

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

@NeedHelp__ You sure? I'm assuming there was a 0 since you have 4 numbers in your options.

OpenStudy (anonymous):

yeah there was but I need help with another problem if you could help :)

OpenStudy (anonymous):

Evaluate f(x) = -x2 - 4 for x = -3 a. -9 b. -13 c. 1 d. -1

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

f(x) = -1[(-3)(-3)] - 4

OpenStudy (anonymous):

solve what's in the parenthesis ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wouldn't that cancel out?

OpenStudy (anonymous):

Nope, (-3)(-3) = 9 so: f(x) = -1[9] - 4

OpenStudy (anonymous):

a negative multiplied by a negative = a positive.

OpenStudy (anonymous):

oh yeah >.< okay

OpenStudy (anonymous):

Since exponents is just a means of saying how many times a number multiply by itself, all even exponents of a negative number will give a positive. All odd exponents will give a negative number.

OpenStudy (anonymous):

\[-x ^{2} = x; -x ^{3} = -x\]

OpenStudy (anonymous):

So what's the answer?

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

Go on: f(x) = -1[9] - 4

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

f(x) = -9-4

OpenStudy (anonymous):

Oh.. >.< -13

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

evaluate F(x) = -2x - 5 for x = 3 a. -11 b. 11 c. 1 d. -6

OpenStudy (anonymous):

same as before :) Replace x with 3 then PEMDAS

OpenStudy (anonymous):

I think its -6 but im sure im wrong

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

I mean if you just mean -2x then yes

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

f(x) = -6-5

OpenStudy (anonymous):

oh -11

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

evaluate f (x) = 1/4x for x = -5 A. -4/5 B. -1 1/4 C. -1/20 D. 20

OpenStudy (anonymous):

Does the equation look like so: \[f(x) = \frac{ 1 }{ 4x }?\]

OpenStudy (anonymous):

yes the x is separate though

OpenStudy (anonymous):

Like this: \[f(x) = \frac{ 1 }{ 4 }x\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Multiply x with 1/4.

OpenStudy (anonymous):

Basically do this: Since all whole numbers, such as 1, 2, 3, are simply \[\frac{ a }{ 1 }\] then when multiplying a whole number with a fraction the equation would look like so:\[\frac{ 1 }{ 4 } (\frac{ x }{ 1 }) = \frac{ 1 }{ 4 } (\frac{ -5 }{ 1 }) = \frac{ 1(-5) }{ 4(1) }\]

OpenStudy (anonymous):

-4/5?

OpenStudy (anonymous):

You got them backwards.

OpenStudy (anonymous):

-5/4 sorry

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

but that's not an option

OpenStudy (anonymous):

it is :)

OpenStudy (anonymous):

no the only one that has those numbers is A which is -4/5 >.<

OpenStudy (anonymous):

Improper fractions are rational numbers where the numerator is greater than the denominator. Improper fractions can be rewritten as a mixed number –an integer plus a proper fraction. An improper fraction represents a number greater than one.

OpenStudy (anonymous):

20??

OpenStudy (anonymous):

"/" This donates division.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

5 divided by 4?

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

0.25 in fractions?

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } = 0.5\]

OpenStudy (anonymous):

1 divided by 2 = 0.5

OpenStudy (anonymous):

1 divided by 4 = 0.25

OpenStudy (anonymous):

Answer: \[-1\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

25/100

OpenStudy (anonymous):

25/100 divide both by 25 so that the numerator is 1.

OpenStudy (anonymous):

You'll get 1/4

OpenStudy (anonymous):

Simpler would have been to divide 5 with 4 and keep the remainder as a fraction.

OpenStudy (anonymous):

Example: \[\frac{ 34 }{ 7 } = 4 \frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

Since 34/7 = 4 r 6

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

yes thanks this is the last one c: f(x) = -4x - 5 for x = -1 A. -1 B. -9 C. 1 D. 4

OpenStudy (anonymous):

It's the same as before. I believe, you can try this on your own.

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@Sadworld

OpenStudy (anonymous):

Try it again. Show me your steps, so I can show you were you got it wrong.

OpenStudy (anonymous):

i see what i did wrong now i multiplied where i wasn't suppose to

OpenStudy (anonymous):

@NeedHelp__ yes :)

OpenStudy (anonymous):

\[f(-1)=-4(-1)-5\]

OpenStudy (anonymous):

Actually you added where you suppose to not.

OpenStudy (anonymous):

so wat u get? @NeedHelp__

OpenStudy (anonymous):

two negatives give a positive.

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

4-5?

OpenStudy (anonymous):

=-1(-4)-5 =4-5 =? solve it

OpenStudy (anonymous):

Previously, you had the same problem. It's the same as before.

OpenStudy (anonymous):

I quote: "f(x) = -x + 4"

OpenStudy (anonymous):

so it would be negative one

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

thanks c: